1.     Determine the critical points of each of 




1
Expert's answer
2021-12-23T16:04:10-0500

for example:

"f(x) = x^2e^x"


then:

"f'(x) =2xe^x+ x^2e^x=x(2+x)e^x=0"


at critical point "x=0" :

f'(x) changes its sign from "-" to "+", so this is local minimum


at critical point "x=-2" :

f'(x) changes its sign from "+" to "-", so this is local maximum